OpenStudy (anonymous):

I would really appreciate help finding the concavity for this function, i can't seem to get it. f(x) = { x^3 }/{ x^2 - 25 }

OpenStudy (anonymous):

i can see why because the second derivative sucks to find. but once you have it, it is easy.

OpenStudy (anonymous):

ok i lied, because i did it wrong. it is still a pain

OpenStudy (anonymous):

\[f''(x)=\frac{50x(x^2+75)}{(x^2-25)^3}\]

OpenStudy (anonymous):

\[x^2+75>0\] so we can ignore that part \[(x^2-25)^3> 0\iff x^2-25>0\] so we can forget about the cube. we only need to consider this \[\frac{x}{x^2-25}=\frac{x}{(x+5)(x-5)}\] which will change sign at \[-5,0,5\] so we need to consider the intervals \[(-\infty, -5),(-5,0),(0,5), (5,\infty)\]

OpenStudy (anonymous):

it will alternate sign over these intervals, and you can check that it is positive on \[(-5,0)\cup (5,\infty)\] and negative on the other two. so those are the intervals over which your function is concave up and it will be concave down on \[(-\infty,-5)\cup (0,5)\]

OpenStudy (anonymous):

I have to mention too that its defined on the interval (-20,18)

OpenStudy (anonymous):

does that make any difference?

OpenStudy (anonymous):

then adjust accordingly which just means change minus infinity to -20 and infinity to 18

OpenStudy (anonymous):

so the concavity would be up on -5,0 and also on 5,18 and then it would be down on -20,-5 and 0,5?

OpenStudy (anonymous):

I got it thank you so much!